Question 44
ProbabilityA bag contains 4 white balls and 6 red balls. Two balls are taken from the bag without replacement. What is the probability that they are both red?
- A.1/3
- B.2/9
- C.2/15
- D.1/5
- E.3/5.
Show the answer and explanation
Answer: A (1/3)
Explanation
P(first red) = 6/10 and P(second red) = 5/9, so P(both red) = 6/10 × 5/9 = 1/3.